OpenStudy (anonymous):

f '(t)=t + (1/t^3), t>0, f(1)=6 Find f I got f (t) = (t^2/2) + (1/2t^2) + c .. stuck after that..

OpenStudy (anonymous):

oh! i put the one in where t's are, and make it equal to 6..?

OpenStudy (anonymous):

i think you integrated \(\frac{1}{t^3}\) wrongly

OpenStudy (anonymous):

i did?

OpenStudy (anonymous):

oh it should be -2

OpenStudy (anonymous):

yeah looks like so \[\frac{1}{t^3} => t^{-3} => \int t^{-3}dt\]

OpenStudy (anonymous):

it shoulkd be \(\frac{1}{2x^2}\)

OpenStudy (anonymous):

there's a minus infront of it lol

OpenStudy (anonymous):

yes so f (t) = (t^2/2) - (1/2t^2) + c

OpenStudy (anonymous):

yes, sub in 1 then set it = 6 then calculate for c then you're done

OpenStudy (anonymous):

like t= 1

OpenStudy (anonymous):

c= 5! thanks for the catch!

OpenStudy (anonymous):

:)